Show that the geometric mean between x and y is ±√xy and the common ratio r is r=^n+1√y/a
1
Expert's answer
2020-06-22T18:16:21-0400

If three quantities are in Geometric Progression then the middle one is called the geometric mean of the other two.

Let, three numbers "x, G" and "y" are in Geometric Progression then, the middle number "G" is called the geometric mean between two numbers "x" and "y."

"x, G" and "y" are in Geometric Progression "G\not=0, xy>0"


"<=>\dfrac{G}{x}=\dfrac{y}{G}"


"<=>xy=G^2"


"G=\pm\sqrt{xy}"



Insert "n" geometric means between "a" and "y."

Let "a_1, a_2, ..., a_n" be "n" geometric means between "a" and "y."

The numbers "a,a_1, a_2, ..., a_n, y" are in Geometric Progression.

Common ratio


"r={a_1\over a}"

Then


"y=ar^{n+1}"

Hence


"r=\sqrt[n+1]{{y\over a}}"